OpenStudy (pphalke):

Sam has five favorite football teams, and every week, he puts their flags on his flagpole in random order a. in a given week, how many different ways can he arrange the flags? b. If he only has room for three of the flags on his flagpole, how many different ways can they be arranged?

OpenStudy (pphalke):

@welshfella please help do u know this?

OpenStudy (pphalke):

@princeharryyy @jigglypuff314 @jhonyy9

OpenStudy (princeharryyy):

5*5*5*5*5*5*5 for the first one

OpenStudy (princeharryyy):

3! = 3*2*1= 6 ways for second.

OpenStudy (pphalke):

the answer for b is 60 ways

OpenStudy (princeharryyy):

hold on let me check it again.

OpenStudy (princeharryyy):

sorry. IT's 5*4*3

OpenStudy (princeharryyy):

first flag in 5 ways second in 4 and third in 3 ways.

OpenStudy (princeharryyy):

as he can choose one out of five flag for first one. now one is used he is left with 4 he can use any of 4 so 4 ways and for the last flag he has 3 left he can use any of three so 3 ways. So 5*4*3 = 60

OpenStudy (princeharryyy):

At first I though he had only 3 flags for second question so i wrote 3! but he had 5 flags in the statement before the question. So Don't mind it please. @pphalke